# Learning in Deep Factor Graphs with Gaussian Belief Propagation

### v1

@@ Abstract
We propose an approach to do learning in Gaussian factor graphs. We treat all relevant quantities (inputs, outputs, parameters, latents) as random variables in a graphical model, and view both training and prediction as inference problems with different observed nodes. Our experiments show that these problems can be efficiently solved with belief propagation (BP), whose updates are inherently local, presenting exciting opportunities for distributed and asynchronous training. Our approach can be scaled to deep networks and provides a natural means to do continual learning: use the BP-estimated parameter marginals of the current task as parameter priors for the next. On a video denoising task we demonstrate the benefit of learnable parameters over a classical factor graph approach and we show encouraging performance of deep factor graphs for continual image classification on MNIST.
@@ 1 Introduction
Deep learning (DL) has been shown to be effective for solving many scene understanding tasks. However, neural networks (NNs) remain a poor fit in applications where we require efficient, robust representations which can be trained incrementally. An agent with an accurate and continually updating “world model” of its surroundings, including semantic and compositional understanding of objects, is not limited to reactive behaviour, but can attempt novel reasoning via mental simulation  (Stoica:etal:TechReport2017).
Such a world model will probably, at least in the near term, contain various human-created representations and structures, such as an explicit 3D map of a scene. However, it will also need to be able to learn its own abstract representations of complex entities such as objects or movement patterns. The following related properties of its learning mechanism are important:
@@
    The primary mode of learning should be un- or self-supervised, meaning that a running system can learn useful abstractions without external help.
  2. 2.
    Learning should be continual, taking place incrementally within a running system whose performance is gradually improving in consequence.
  3. 3.
    Learning should run seamlessly alongside existing tried-and-tested hand-designed inference algorithms.
To satisfy desiderata 2 and 3, an agent must be able to fuse multiple models, balancing old vs new, hand-crafted vs learnt. Bayesian inference provides a principle by which multiple signals can be fused: they can be combined according to the rules of probability. The integration of standard DL with incremental probabilistic inference on graphs has so far been with limited success, due to the lack of such principles.
Concurrently, the ability to distribute model training over multiple processing nodes is becoming increasingly important with the growth of i) larger models which must be spread over multiple processors, ii) processors whose cores comprise significant local memory (Graphcore), iii) parallel, distributed and heterogeneous embedded devices (Sutter:Jungle2011). Efficient model-parallel training of NNs with backpropagation is currently limited by backward locking: processors for earlier layers are idle when waiting for the backward error signal. Again, we argue that Bayesian principles provide an answer here: local signals can be fused according to the rules of probability, and this enables effective communication between different regions of a model without relying on a global loss.
Figure 1: In GBP Learning, we design factor graphs whose structure mirrors common neural network architectures, but with learnable parameters included as variables in the graph, and non-linear factors (which have soft-switching behaviour) akin to neurons. The example architecture shown here for image classification comprises a convolutional layer, a max pooling layer and a dense projection layer. A common set of blue parameter variables (circles) is shared across all image/label observations, while there are copies of the red activation and green input/output variables for each observation. Factors (black squares) between layers constrain their representations to be locally consistent, while those attached to inputs and outputs encourage consistency with observations. This example shows a supervised learning setup, but the same architecture could be trained without label supervision by removing the output observation factor. Given such a factor graph, we can train and make predictions in a distributed manner with GBP.
In this paper we propose a new Bayesian approach to machine learning, where parameter learning and communication between layers is part of the same joint procedure. We propose that there is no fundamental difference between traditional “state estimation” via probabilistic inference, and model learning in Bayesian ML. We believe our joint approach to inference and learning distinguishes it from existing methods of Bayesian DL which view them as separate procedures.
Much recent work has shown that Gaussian Belief Propagation (GBP) permits convergent, efficient, incremental inference across a broad range of probabilistic graphical models, including models with non-linear factors and robust cost functions  (Davison:Ortiz:ARXIV2019; Ortiz:etal:CVPR2020; Murai:etal:ARXIV2022; Patwardhan:etal:ARXIV2022). Our hypothesis is that this machinery can be applied to general machine learning, which will take place via inference on large, overparameterised graphs with non-linear Gaussian factors, and architectures resembling common NNs. We call this new approach GBP Learning. We claim that the important features of DL are i) randomly initialised, massively over-parameterised networks (allen2019learning) with ii) architectural motifs which encode good inductive bias, and iii) non-linearities which switch to selectively activate and prune when exposed to data (glorot2011deep) — rather than the specifics of the commonly used backpropagation learning rule. We believe that we can embed these properties in Gaussian factor graphs, while gaining the advantages of probabilistic modelling for continual learning and interfacing with other Bayesian components. Such models can be trained with only local updates using GBP.
Our paper proceeds as follows. First, we briefly describe GBP in Section 2, before introducing our model and how it is trained in Section 3. Experimental results are given in Section 4, before summarising previous work (Section 5) and concluding (Section 6). In short, our contributions are:
  1. 1.
    A new class of deep networks which can be represented as Gaussian factor graphs and trained with BP.
@@ 2.1 Factor graphs
Here $\mathbf{x}_{\phi_{j}}$ denotes the vector of all $V_{j}$ variables in the neighbourhood of $\phi_{j}$ and $Z$ is a normalising constant. A factor graph is bipartite: variables only connect to factors and vice versa. Each factor may encode an observation, or prior about a variable or group of variables. Factor functions $\phi_{j}(\cdot)$ may unnormalised distributions and relate to their energy, $E_{j}(\cdot)$, as
@@
$E_{j}\left(\mathbf{x}_{\phi_{j}}\right)=\frac{1}{2}\left(\mathbf{y}_{j}-\mathbf{h}\left(\mathbf{x}_{\phi_{j}}\right)\right)^{{}^{\top}}{\Lambda_{\mathbf{y}_{j}}}\left(\mathbf{y}_{j}-\mathbf{h}\left(\mathbf{x}_{\phi_{j}}\right)\right)~{},$ (3)
@@ 2.2 Belief Propagation
Belief Propagation (pearl1988probabilistic) is a message passing algorithm which can perform inference on factor graphs via distributed, iterative computation. Each message is passed along the edge between one factor $\phi_{j}$ and one variable $x_{i}$. Messages travel in both directions, and we use the notations $m_{{\phi_{j}\rightarrow x_{i}}}$ and $m_{{x_{i}\rightarrow\phi_{j}}}$ for the two types of message. A message takes the form of a probability distribution in the space of the variable involved. The following update rules are iterated until convergence:
@@
$m_{{\phi_{j}\rightarrow x_{i}}}\left(x_{i}\right)$$\leftarrow\sum_{\{x_{n}\}_{\mathrm{ne}\left(j\right)\setminus i}}\phi_{j}\left(\mathbf{x}_{\phi_{j}}\right)\prod_{n\in\mathrm{ne}\left(j\right)\setminus i}m_{{x_{n}\rightarrow\phi_{i}}}\left(x_{n}\right),$ (5)
where $\mathrm{ne}\left(i\right)\setminus j$ indicates the indices of all nodes connected to node $i$, except $j$. After convergence the posterior marginal of a variable can be estimated by the product of its incoming messages:
@@
where $Z_{i}$ is straightforward to compute as $p(x_{i})$ is assumed to belong to a known parametric family. In this work, we assume the variable nodes are univariate, but BP can be extended for multivariate posterior inference by passing vector messages between sets of variables and factors. BP updates (4), (5), (6) have a number of interesting properties. First, all the required computations rely only on the local state of the graph; no global context is necessary. Second, by working in terms of natural parameters, taking the product of a set of messages is reduced to the addition of their parameters.
@@ 2.2.1 Gaussian Belief Propagation
In Gaussian factor graphs, messages $m_{{x_{i}\rightarrow\phi_{j}}}\left(x_{i}\right)$ and $m_{{\phi_{j}\rightarrow x_{i}}}\left(x_{i}\right)$ are normal distributions, parameterised by precision (inverse covariance) matrix $\Lambda$ and information vector $\eta$. The products in the variable to factor message formula (4) become sums:
@@
As in the general case, GBP is not guaranteed to converge in graphs with cycles. However, if it does converge in linear-Gaussian models, it is guaranteed to converge to the correct posterior means(weiss1999correctness). In practice, it has been empirically shown that this correct convergence reliably happens in many types of factor graph. We refer the reader to ortiz2021visual for an in-depth introduction to GBP.
@@ 2.3 Non-linear Factors
GBP supports the use of factors which include non-linear transformations of their connected variables, and this is crucial in our proposed use here for representation learning. Every time we compute messages from a non-linear factor, we linearise it around the current estimates of the variables, $\mathbf{X}_{0}$, i.e. for each factor $\mathbf{h}\left(\mathbf{x}_{\phi_{j}}\right)\approx\mathbf{h}\left(\mathbf{x}_{\phi_{j},0}\right)+\mathbf{J}_{j}^{{}^{\top}}\left(\mathbf{x}_{\phi_{j}}-\mathbf{x}_{\phi_{j},0}\right)$ where $\mathbf{J}_{j}:=\left.\frac{\partial\mathbf{h}\left(\mathbf{x}_{\phi_{j}}\right)}{\partial\mathbf{x}_{\phi_{j}}}\right|_{\mathbf{x}_{\phi_{j},0}}$. Linearisation can be implemented using the following forms of factor precision and information in (8):
@@
See Section 3.3 of Davison:Ortiz:ARXIV2019 for details and derivation.
@@ 3 GBP Learning
Our aim is to produce factor graphs which share similar architectural inductive biases to NNs, that can be overparameterised in a similar way, but can be trained with GBP. We will now describe the key factors in our model and our efficient GBP routine for training and prediction. Throughout this section we will use “inputs” to mean the variables within a layer which are closest to the pixels (in terms of the number of edges along the shortest path). Conversely “outputs” is used to mean the variables in the layer which are furthest from the pixels. However, BP is bidirectional so these quantities should not be thought of as inputs or outputs in the sense of a function.
@@ 3.1 Deep Factor Graphs
Our networks are designed to find representations based on local consistency. A layer $l$ in our graph learns to produce a representation $\mathbf{x}_{l}\in\mathbb{R}^{D_{l}}$ that generates the previous layer $\mathbf{x}_{l-1}\in\mathbb{R}^{D_{l-1}}$ or subsequent layer $\mathbf{x}_{l-1}\in\mathbb{R}^{D_{l+1}}$, via a non-linear transformation. Applying this principle in e.g. the generative direction, together with the Gaussian assumption, suggests the following form for the factor energy:
@@
which is low when $\mathbf{f}(\mathbf{x}_{l},\mathbf{\Theta}_{l})$ matches the input $\mathbf{x}_{l-1}$ using parameters $\mathbf{\Theta}_{l}$ and output $\mathbf{x}_{l}$. $\sigma_{l}$ is the factor strength ($\Lambda_{\mathbf{y}_{l}}=\frac{1}{\sigma_{l}^{2}}\mathbf{I}_{D_{l-1}}$). Through choice of $\mathbf{f}(\cdot,\cdot)$ we can encode different operations and inductive biases.
CNNs, which have been successfully applied across a range of computer vision tasks, have a sparse connectivity structure which suggests efficient factor analogues. In our convolutional layers, a factor at spatial location $(a,b)$ connects to i) the $K_{l}\times K_{l}$ patch of the input within its receptive field, $\mathbf{X}_{l-1}^{(a,b)}\in\mathbb{R}^{K_{l}\times K_{l}\times C_{l-1}}$, ii) the corresponding activation variable for an output channel $c$, $\mathbf{x}_{l}^{(a,b,c)}\in\mathbb{R}^{C_{l}}$ and iii) the parameters: filters $\mathbf{\theta}_{l}^{(c)}\in\mathbb{R}^{K_{l}\times K_{l}\times C_{l-1}}$ and bias $b_{l}^{(c)}$ for output channel $c$, which are shared across the layer. The energy can be written as:
@@
We can similarly define a transposed convolution layer. In this case, each filter is weighted by the output variable of its corresponding channel, and the weighted sum reconstructs the inputs. Note that for a stride smaller than the kernel size, each input will belong to multiple receptive fields and they are summed over to give its reconstruction. For example, for a stride of one, and neglecting edge effects we have:
$E_{\mathrm{convT}}^{(a,b,c)}$ $=\frac{\left(x_{l-1}^{(a,b,c)}-r\left(\mathbf{X}_{l}^{(a-i,b-j)},\mathbf{\theta}_{l}^{(i,j,c)},b_{l}^{(c)}\right)\right)^{2}}{2\sigma_{l}^{2}}~{}.$ (13)
@@
In addition to convolutions, we introduce factors akin to other common CNN layers. For example, we use max-pooling factors to reduce the spatial extent of the representation. For an input patch from channel $c$, $\mathbf{X}_{l-1}^{(a,b,c)}\in\mathbb{R}^{K\times K}$, centred at $(a,b)$, the connected pooling factor has energy:
@@
$E_{\mathrm{dense}}$ $=\frac{\left\lVert\mathbf{x}_{L}-\left(\mathbf{W}_{L}^{\top}\mathrm{vec}\left(\mathbf{X}_{L-1}\right)+\mathbf{d}_{L}\right)\right\rVert_{2}^{2}}{2\sigma_{L}^{2}}~{}.$ (15)
Last, class supervision may then be incorporated by treating last layer activations $\mathbf{x}_{L}$ as logits and connecting them to a observation factor with energy:
$E_{\mathrm{softmax}}$ $=\frac{\left\lVert\mathrm{softmax}\left(\mathbf{x}_{L}\right)-\mathbbm{1}_{y}\right\rVert_{2}^{2}}{\sigma_{\mathrm{softmax}}^{2}}~{},$ (16)
where $\mathbbm{1}_{y}$ is a one-hot encoding of the observed class $y$.
@@
The “layer” abstractions defined above can be composed to produce deep models with similar design freedom to NNs. Similarly, our models may be overparameterised by introducing large numbers of learnable weights and we believe that the inclusion of non-linear activation functions $\mathbf{g}(\cdot)$ in our factor graph can aid representation learning as in DL. In particular, we note that for non-linear factors the Jacobian is a function of the variable values $\mathbf{J}_{j}=\mathbf{J}_{j}\left(\mathbf{x}_{\phi_{j},0}\right)$, which will cause the strength of the factors (9) to vary depending on the input data. We expect this to produce a similar “soft switching” of connections as observed in non-linear NNs. Note that the set of layers described here are non-exhaustive, and the translation of other neural architectures into factor graphs may be a worthwhile future direction.
3.2 Continual Learning

### v2

@@ Abstract
We propose an approach to do learning in Gaussian factor graphs. We treat all relevant quantities (inputs, outputs, parameters, activations) as random variables in a graphical model, and view training and prediction as inference problems with different observed nodes. Our experiments show that these problems can be efficiently solved with belief propagation (BP), whose updates are inherently local, presenting exciting opportunities for distributed and asynchronous training. Our approach can be scaled to deep networks and provides a natural means to do continual learning: use the BP-estimated posterior of the current task as a prior for the next. On a video denoising task we demonstrate the benefit of learnable parameters over a classical factor graph approach and we show encouraging performance of deep factor graphs for continual image classification.
@@ 1 Introduction
Figure 1: In GBP Learning, we design factor graphs whose structure mirrors common NN architectures, enabling distributed training and prediction with GBP. Learnable parameters are included as random variables (circles), as are inputs, outputs and activations. The parameters are shared over across all observations, where the other variables are copied once per observation. Factors (black squares) between layers constrain their representations to be locally consistent, while those attached to inputs and outputs encourage compatibility with observation. Non-linear inter-layer factors to enable soft-switching behaviour. This example architecture for image classification comprises convolutional, max pooling and dense projection layers. The same architecture could learn in an unsupervised way by removing the output observation factor.
Deep learning (DL) has been transformative across many domains. However, its applicability is limited in cases where i) we require efficient, robust representations which can be trained incrementally; ii) supervision is sparse or irregular; and iii) learning must augment, or run alongside, hand-designed solvers. Pretrained models, when available, might mitigate these limitations, but struggle as train and test distributions diverge. How they should be updated online remains an open research question.
Concurrently, we reflect on how neural networks (NNs) are trained. Despite backpropagation rumelhart1985learning being largely successful for DL, training NNs spread over multiple processors is made difficult by backward locking: processors for earlier layers sit idle after their part in the forward pass, awaiting the backward error signal. This challenge will become more pertinent with the growth of i) larger models which must be distributed over many devices, ii) new hardware architectures whose cores have significant local memory (Graphcore; cerebras), and iii) parallel, distributed and heterogeneous embedded devices (Sutter:Jungle2011). We thus expect a growing need for more flexible training algorithms which admit efficient model-parallelism.
We argue the above challenges essentially relate to the fusion of multiple signals: old and new (for incremental learning); hand-crafted vs learnt; and representations between different layers of a model (distributed training). Bayesian principles offer a clear answer on how to fuse: signals should be combined according to the rules of probability. We seek to exploit this fact in our probabilistic approach to DL. Specifically, our models are factor graphs (Fig. 1) with random variables for all quantities relevant to DL: inputs, outputs, activations and parameters. This representation enables continual learning via online updating of the parameter posterior, and interoperability through connections with other factor graphs. We seek to design the models so as to retain the properties which we believe make DL powerful. Namely random initialisation and over-parameterisation (allen2019learning); architectural motifs encoding good inductive biases; and non-linearities which switch to selectively activate and prune when exposed to data (glorot2011deep).
Much recent work has shown that Gaussian BP (GBP) to be a robust and effective algorithm for distributed inference in factor graphs, even in the presence of non-linear and non-Gaussian factors (Davison:Ortiz:ARXIV2019; Ortiz:etal:CVPR2020; Murai:etal:ARXIV2022; Patwardhan:etal:ARXIV2022). It is thus our choice of inference engine here. By approximating the factors in our model as Gaussian, we can use of GBP for training (inference over parameters, given observations) and prediction (inference of e.g. outputs, given parameters and inputs). The generality of GBP provides flexibility as to which variables are observed, meaning training and prediction are essentially the same computation, and partial observations or missing labels do not require fundamentally different treatment. As BP is inherently local and stateful, our training can be distributed and asynchronous.
Our work shares similar goals to lucibello2022deep(lucibello2022deep), who train MLP-like factor graphs with GBP using analytically derived message updates. However, their approach lacks generality: the derived updates hold only for dense layers, and they focus on models with binary weights and sign activation functions. Here, in contrast, we present a general method to train deep models with GBP, requiring only that appropriate energy functions be specified for inter-layer factors, and naturally supporting continuous weights. Message updates need not be rederived for different architectures.
We call our approach GBP Learning. In short, we construct factor graph architectures inspired by those in DL (Section 3.1). Within these models, GBP inference admits flexible and distributed training and prediction (Section 3.2) and we can learn incrementally via Bayesian filtering over parameters (Section 3.4). Our experiments demonstrate the benefit of factor graphs with learnable parameters over hand-designed solvers in a video denoising task (Section 5.2). For image classification (Section 5.3), we evaluate our continual learning approach by single epoch training on MNIST. We achieve performance equivalent to an Adam-trained CNN with a replay buffer of $6\times 10^{3}$ examples, and show this performance is robust to asynchronous training. Last, we compare to lucibello2022deep(lucibello2022deep), outperforming them on both MNIST (by $0.8\%$) and CIFAR10 (by $11.8\%$).
Our main contributions are as follows:
@@
    A new class of deep networks which we represent as Gaussian factor graphs and train with GBP.
@@ 2.1 Factor graphs
Here $\mathbf{x}_{\phi_{j}}$ denotes the vector of all $V_{j}$ variables in the neighbourhood of $\phi_{j}$ and $Z$ is a normalising constant. A factor graph is bipartite: variables only connect to factors and vice versa. Each factor may encode an observation, or prior about one of many variables. Factor functions $\phi_{j}(\cdot)$ may be unnormalised distributions relating to their energy, $E_{j}(\cdot)$, as
@@
$E_{j}\left(\mathbf{x}_{\phi_{j}}\right)=\frac{1}{2}\left(\mathbf{y}_{j}-\mathbf{h}\left(\mathbf{x}_{\phi_{j}}\right)\right)^{{}^{\top}}{\Lambda_{\mathbf{y}_{j}}}\left(\mathbf{y}_{j}-\mathbf{h}\left(\mathbf{x}_{\phi_{j}}\right)\right)$ (3)
@@ 2.2 Belief Propagation
Belief Propagation (pearl1988probabilistic) is a message passing algorithm which can perform inference on factor graphs via distributed, iterative computation. Each message is passed along the edge between one factor $\phi_{j}$ and one variable $x_{i}$. Messages travel in both directions, and we use the notations $m_{{\phi_{j}\rightarrow x_{i}}}$ and $m_{{x_{i}\rightarrow\phi_{j}}}$ for the two types of message. Each message is a probability distribution in the space of the variable involved. The following update rules are iterated until convergence:
@@
$m_{{\phi_{j}\rightarrow x_{i}}}\left(x_{i}\right)$$\leftarrow\!\!\!\!\!\!\sum_{\{x_{n}\}_{\mathrm{ne}\left(j\right)\setminus i}}\!\!\!\!\!\phi_{j}\left(\mathbf{x}_{\phi_{j}}\right)\!\!\!\!\!\prod_{n\in\mathrm{ne}\left(j\right)\setminus i}\!\!\!\!\!m_{{x_{n}\rightarrow\phi_{i}}}\left(x_{n}\right)~{},$ (5)
where $\mathrm{ne}\left(i\right)\setminus j$ is the indices of all nodes connected to $i$, except $j$. After convergence the posterior marginal of a variable is estimated by the product of its incoming messages:
@@
where $Z_{i}$ is straightforward to compute if the messages belong to a known parametric family. In this work, we assume the variable nodes are univariate, but BP can be extended for multivariate posterior inference by passing vector messages between sets of variables and factors. BP updates (4), (5), (6) have a number of interesting properties. First, the required computations rely only on the local state of the graph; no global context is necessary. Second, for exponential family distributions under natural parameterisation, taking the product of a set of messages is reduced to the addition of their parameters.
@@ 2.2.1 Gaussian Belief Propagation
In Gaussian factor graphs, messages $m_{{x_{i}\rightarrow\phi_{j}}}\left(x_{i}\right)$ and $m_{{\phi_{j}\rightarrow x_{i}}}\left(x_{i}\right)$ are normal distributions, with natural parameters being the precision (inverse covariance) matrix $\Lambda$ and information vector $\eta$. The products in the variable to factor message formula (4) become sums:
@@
As in the general case, GBP is not guaranteed to converge in graphs with cycles. However, if it does converge in linear-Gaussian models, it is guaranteed to converge to the correct posterior means (weiss1999correctness). In practice, it has been empirically shown that it converges to correct means in in many types of factor graph. We refer the reader to ortiz2021visual for an in-depth introduction to GBP.
@@ 2.3 Non-linear Factors
GBP supports the use of factors which include non-linear transformations of their connected variables, and this is crucial in our proposed use here for representation learning. Every time a message from a non-linear factor is computed, the factor is linearised around the current estimates of the variables, $\mathbf{X}_{0}$, i.e. for each factor $\mathbf{h}\left(\mathbf{x}_{\phi_{j}}\right)\approx\mathbf{h}\left(\mathbf{x}_{\phi_{j},0}\right)+\mathbf{J}_{j}^{{}^{\top}}\left(\mathbf{x}_{\phi_{j}}-\mathbf{x}_{\phi_{j},0}\right)$ where $\mathbf{J}_{j}:=\left.\partial\mathbf{h}\left(\mathbf{x}_{\phi_{j}}\right)/\partial\mathbf{x}_{\phi_{j}}\right|_{\mathbf{x}_{\phi_{j},0}}$. Linearisation can be implemented using the following forms of factor precision and information in (8):
@@
See Section 3.3 of Davison:Ortiz:ARXIV2019 for details.
@@ 3 GBP Learning
Our aim is to produce factor graphs which have similar architectural inductive biases to NNs, that can be overparameterised in a similar way, but can be trained with GBP. We will now describe the key factors in our model and our efficient GBP routine for training and prediction. The factor energies of form (3) contain $\mathbf{h}\left(\cdot\right)$ and $\mathbf{y}$ which, along with $\Lambda_{\mathbf{y}}$, are sufficient to deduce the linearised factor (9). This, in turn, defines the GBP message updates (7), (8). It is thus sufficient to describe our model in terms of factor energies.
@@ 3.1 Deep Factor Graphs
We design networks to find representations based on local consistency, i.e. the activations $\mathbf{x}_{l}\in\mathbb{R}^{D_{l}}$ in a layer $l$ should “predict” those in either the previous layer $\mathbf{x}_{l-1}\in\mathbb{R}^{D_{l-1}}$ or subsequent layer $\mathbf{x}_{l+1}\in\mathbb{R}^{D_{l+1}}$, via a parametric non-linear transformation $\mathbf{f}(\cdot,\mathbf{\Theta}_{l})$. Applying this principle in e.g. the generative direction, together with the Gaussian assumption, suggests the following form for the factor energy:
@@
which is low when $\mathbf{f}(\mathbf{x}_{l},\mathbf{\Theta}_{l})$ matches the input [^1] We use “inputs” to mean the activation variables within a layer which are closer to the pixels, and “outputs” those which are further away. However, BP is bidirectional so these quantities are not equivalent to the inputs of a function. $\mathbf{x}_{l-1}$ using parameters $\mathbf{\Theta}_{l}$ and output $\mathbf{x}_{l}$. $\sigma_{l}$ is the factor strength ($\Lambda_{\mathbf{y}_{l}}=\frac{1}{\sigma_{l}^{2}}\mathbf{I}_{D_{l-1}}$). Through choice of $\mathbf{f}(\cdot,\cdot)$ we can encode different operations and inductive biases.
CNNs, which have been successfully applied across a range of computer vision tasks, have a sparse connectivity structure which suggests efficient factor graph analogues. In our convolutional layers, a factor at spatial location $(a,b)$ connects to i) the $K_{l}\times K_{l}$ patch of the input within its receptive field, $\mathbf{X}_{l-1}^{(a,b)}\in\mathbb{R}^{K_{l}\times K_{l}\times C_{l-1}}$, ii) the corresponding activation variable for an output channel $c$, $\mathbf{x}_{l}^{(a,b,c)}\in\mathbb{R}^{C_{l}}$ and iii) the parameters: filters $\mathbf{\theta}_{l}^{(c)}\in\mathbb{R}^{K_{l}\times K_{l}\times C_{l-1}}$ and bias $b_{l}^{(c)}$ for output channel $c$, which are shared across the layer. The energy can be written as:
@@
We can similarly define a transposed convolution layer. In this case, each filter is weighted by the output variable of its corresponding channel, and the weighted sum reconstructs the inputs. Note that for a stride smaller than the kernel size, each input will belong to multiple receptive fields and they are summed over to give its reconstruction. For example, for a stride of one and neglecting edge effects, these contributions can be combined according to:
$E_{\mathrm{convT}}^{(a,b,c)}$ $=\frac{\left(x_{l-1}^{(a,b,c)}-r\left(\mathbf{X}_{l}^{(a,b)},\mathbf{\theta}_{l}^{(c)},b_{l}^{(c)}\right)\right)^{2}}{2\sigma_{l}^{2}}~{}.$
@@
$E_{\mathrm{dense}}=\frac{\left\lVert\mathbf{x}_{L}-g\left(\mathbf{W}_{L}^{\top}\mathrm{vec}\left(\mathbf{X}_{L-1}\right)+\mathbf{d}_{L}\right)\right\rVert_{2}^{2}}{2\sigma_{L}^{2}}~{}.$ (13)
where the activation function $g(\cdot)$ is included for generality, but set to identity for image classification.
In addition to convolutional and dense factors, we design factors akin to other common CNN layers. For example, we use max-pooling factors to reduce the spatial extent of the representation, upsampling layers to increase it and softmax observation factors for class supervision. We include the energies for these factors in App. A. Note the the set of layers described here is non-exhaustive, we leave the exploration of other layer types as future work.
@@
These “layer” abstractions can be composed to produce deep models with similar design freedom to DL. Further, our models may be overparameterised by introducing large numbers of learnable weights and we believe that the inclusion of non-linear activation functions $\mathbf{g}(\cdot)$ in our factor graph can aid representation learning as in DL. In particular, we note that for non-linear factors, the factor Jacobian is a function of the current variable values $\mathbf{J}_{j}=\mathbf{J}_{j}\left(\mathbf{x}_{\phi_{j},0}\right)$, which will cause the strength of the factors (9) to vary depending on the input data. We expect this to produce a similar “soft switching” of connections as observed in non-linear NNs.
3.2 Learning and Predicting with GBP Inference
We have described the components of deep factor graph models, in which inputs, outputs, activations and parameters are random variables. As we have designed these models with (non-linear) Gaussian factors, we can do inference with GBP via iterative linearisation (Section 2.3). We apply this inference engine to estimate posteriors over all latent variables. Note that there is no fundamental difference between training and prediction — only a difference in which variables are observed. In training, we infer a posterior over parameters given observed inputs and, where supervision is available, outputs. To then make predictions on new examples, we run GBP to predict the unobserved inputs/outputs given the observed inputs/outputs and parameters.
Our message schedule proceeds as follows. For a given batch, we initialise the graph and update messages by sweeping forward and backward through the layers: first nearest the input observations, progressing to the deepest layer and back again. We repeat this for a specified number of iterations. Within each layer, we compute all factor to variable updates in parallel, and likewise for the variable to factor messages of each variable type (inputs, outputs, parameters as applicable). In addition, we experimentally show our approach works well with layerwise asynchronous message schedules (Section 5.3.1). While we have demonstrated these schedules work, they are likely suboptimal and we leave exploration for future work. We find that applying damping (murphy2013loopy) and dropout to the factor to variable messages is sufficient for stable GBP.
3.3 Efficient GBP
Efficient inference is necessary for our models to be useful in practice. However, the inversion of a $\left(V_{j}-1\right)\times\left(V_{j}-1\right)$ matrix to compute $\Sigma_{\setminus i,\setminus i}^{\left(\phi_{j}+m\right)}$ in (8) has $O\left((V_{j}-1)^{3}\right)$ complexity, bottlenecking GBP. To alleviate this, we exploit the structure of the matrix being inverted. In particular, i) for factors with observation dimension $M=\dim\left(\mathbf{y}\right)<V_{j}$, the precision $\Lambda^{(\phi_{j})}$ (9) is low-rank, and ii) for graphs with scalar variable nodes, $\mathbf{D}$ is diagonal. Thus their sum may be efficiently inverted via the Woodbury identity (woodbury1950inverting). Further savings come from reusing intermediates when computing messages to multiple variables (9). These optimisations change the complexity of updating messages from a factor to all $V$ variables, from $O(V\left(V-1\right)^{3})$ to $O\left(VM^{3}\right)$. Space complexity is changed from $O(V^{2})$ to $O(VM+M^{2})$. See App. B for details. These results constitute significant savings when $M<<V$ which is usually true in practice.
3.4 Continual Learning and Minibatching